OpenStudy (anonymous):

Find the product. FOR MEDAL PLZ HELP (x^2+5)^2

OpenStudy (lucaz):

you can use (a+b)^2 = a^2+2ab+b^2, in this case a=(x^2) and b=5

OpenStudy (lucaz):

(x^2)^2+2(x^2)(5)+5^2= x^4+10x^2+25